# Timing Matters: Enhancing User Experience through Temporal Prediction in Smart Homes

### v1

@@ Abstract
Have you ever considered the sheer volume of actions we perform using IoT (Internet of Things) devices within our homes, offices, and daily environments? From the mundane act of flicking a light switch to the precise adjustment of room temperatures, we are surrounded by a wealth of data, each representing a glimpse into user behaviour. While existing research has sought to decipher user behaviours from these interactions and their timestamps, a critical dimension still needs to be explored: the timing of these actions. Despite extensive efforts to understand and forecast user behaviours, the temporal dimension of these interactions has received scant attention. However, the timing of actions holds profound implications for user experience, efficiency, and overall satisfaction with intelligent systems. In our paper, we venture into the less-explored realm of human-centric AI by endeavoring to predict user actions and their timing. To achieve this, we contribute a meticulously synthesized dataset comprising 11k sequences of actions paired with their respective date and time stamps. Building upon this dataset, we propose our model, which employs advanced machine learning techniques for k-class classification over time intervals within a day. To the best of our knowledge, this is the first attempt at time prediction for smart homes. We achieve a 40% (96-class) accuracy across all datasets and an 80% (8-class) accuracy on the dataset containing exact timestamps, showcasing the efficacy of our approach in predicting the temporal dynamics of user actions within smart environments.
Introduction
@@
In the era of technological advancement, the rapid growth of Internet of Things (IoT) solutions has resulted in a increase of smart devices within households, which is likely to almost double from 15.1 billion in 2020 to more than 29 billion IoT devices in 2030 (Lionel 2023). This has changed the way we interact with technology in our daily lives. From smart thermostats to voice-activated assistants, these devices have seamlessly integrated into our homes, workplaces, and public spaces, providing convenience, efficiency, and connectivity like never before.
@@
Beyond task automation and real-time data provision, smart devices also have the potential to impact how machines understand and respond to human behaviour. By continuously collecting data on user interactions, preferences, and habits, these devices offer insights into our routines, lifestyle choices, and even our cognitive processes. Leveraging this data can lead to the development of more intuitive and personalized technologies that anticipate the time of our needs and preferences, ultimately enhancing our overall quality of life.
@@
A crucial yet often overlooked aspect of user behavior prediction is the timing of actions. While existing research primarily focuses on predicting the next user action, little attention has been given to forecasting when these actions will occur. Understanding and predicting the timing of user actions is as important as predicting the actions themselves, as it significantly enhances the responsiveness and effectiveness of smart systems.
Consider a smart home system designed to assist with daily routines. If the system can predict that a user will typically start making breakfast at 7:30 AM, it can preheat the oven or prepare the coffee machine just in time, thereby providing a seamless and proactive user experience. Without such temporal predictions, the system remains reactive, merely responding to user commands rather than anticipating needs.
@@
Temporal predictions are integral to the functionality of autonomous systems. Whether developing assistive robots for home use or automating smart home systems, accurate time prediction is crucial for active user assistance. Without temporal prediction, applications remain passive, lacking the capability to proactively support users. By addressing this novel task, our research takes the first steps towards creating automated assistive AI systems capable of significantly improving the user experience in smart homes.
@@
Figure 1: Given the history of actions taken by the user we aim is to predict the time of action
Contributions
@@
    A Synthesized Dataset, which consists of information about 16 devices and 121 device controls. This new dataset has $\sim 12000$ instances each consisting of a sequence of $10$ consecutive actions and the exact timestamps. (Section 5.1)
@@
    A novel approach Timing-Matters for predicting the time of action on open source and our dataset. (Section 4)
@@
    A extensive study of modeling classification problem with the choice of the length of the context window and the window duration. (Section 6)
Related Work
User Behaviour Prediction in Smart Homes
@@
The primary focus of research on user behaviour prediction in smart homes is on learning-based techniques, specifically deep learning and conventional machine learning techniques. Traditional methods like Hidden Markov Models (HMMs)  (Lin, Zheng, and Xiang 2018; Sikder et al. 2019) have been used to extract the user’s behaviour pattern and detect anamolies. However, HMM will fail due to the independence assumption (Eddy 1996) of HMM which makes it unable to consider context information. (Alaghbari et al. 2022; Du, Lim, and Tan 2019; Farayez, Reaz, and Arsad 2018; Tax 2018; Zou et al. 2023) exploit Long Short-Term Memory Networks (LSTM) to predict user behaviour in smart home scenarios. However, LSTM can only model long-term sequential influence (Graves 2012), but misses out on the complex heterogeneous transitions and periodic sequential patterns caused by users’ routines and intents. SmartSense (Jeon et al. 2022) adopts knowledge transfer to exploit user intent and employs a two-stage encoder to mine contextual information. However, it does not classify the routines before inputting the sequence into the encoder. Second, the complex heterogeneous transitions are not modeled. Third, it does not take into account multi-level periodicity due to a lack of consideration of historical sequences. DeepUDI (Xiao et al. 2023a) employs the attention mechanism to consider the periodicity of human behaviour and intent-aware encoder to consider intents. SmartUDI (Xiao et al. 2023b) considers routines, intent, and multi-level periodicity at the same time. However, none of the techniques try to predict the time when the user will perform the action.
Sequential User Behaviour Prediction
@@
Sequential user behaviour prediction methods can be categorized as follows. Traditional methods include Factorizing Personalized Markov Chains (FPMC) (Rendle, Freudenthaler, and Schmidt-Thieme 2010) factorize the users-locations matrix to generate user general preferences to complete the next location prediction. However, the location independence assumption prevents FPMC from capturing complicated sequential information (Rendle, Freudenthaler, and Schmidt-Thieme 2010). CNN-based methods like Caser (Tang and Wang 2018) employs CNN in both the time-axis and feature-axis to capture temporal dynamics in sequential recommendation. However, due to the limited receptive field of CNN, it cannot fully model long-term contextual information (Kim and Kim 2017). RNN-based methods like CARNN (Liu et al. 2016) and SIAR (Rakkappan and Rajan 2019) incorporate contextual information into the RNN for sequential recommendation. However, CARNN and SIAR can not model multi-level periodicity because they only mine contextual information in a single sequence (Liu et al. 2016; Rakkappan and Rajan 2019). Although DeepMove (Feng et al. 2018) uses attention-based RNN to model the correlation between the current sequence and historical sequences, it leads to suboptimal performance and long prediction time since it ambiguously considers all history sequences. SRGNN (Wu et al. 2019; Xu et al. 2019) applies gated GNN to capture complex transition patterns among nodes for a session-based recommendation. However, SRGNN ignores the heterogeneity of the transition patterns caused by users’ intents. SASRec (Kang and McAuley 2018) utilizes unidirectional transformers to capture sequential patterns in sequences while considering the importance of correlations between behaviours. However, SASRec only captures the sequential patterns in single sequences, which prevents it from considering multi-level periodicity.
Problem Statement
This project aims to predict the next time of action given the history of $n-1$ actions taken by the user. Each action comprises the device and the temporal context, like the day of the week, time of day, device, and corresponding action. The goal is to develop a model that takes these input sets and predicts the specific time of action, providing valuable insights into user behaviour. For example, Fig. 1, shows a user sequence of 5 actions and the task of the model is to predict the time of $6^{th}$ action.
TIME OF ACTION PREDICTION
For each instance $u$, we are given
Input
A Sequence of history $S_{u}=[x_{u,1},\ldots,x_{u,(n-1)}]$, where $x_{u,i}=(d_{u,i}^{(1)},d_{u,i}^{(2)},c_{u,i}^{(1)},c_{u,i}^{(2)})$ is a quadruplet of indices of device, device control, day, and time, respectively, of the $i^{th}$ action in the sequence. The entire sequence contains n elements.
Output
The time interval $\hat{y}_{u}$ of the day that the next action will be performed. In our approach, we mainly work with dividing the day into $96$ 15-minute intervals.
@@
Methodology
Time2Vec Embedding
This is a widely used representation of time and ensures that both the sequential and cyclic patterns of time are captured (Kazemi et al. 2019). This representation can be likened to a form of Fourier Transform. Except for the linear element, the rest of the elements split the time into different frequencies. This extracts the cyclic patterns present in the temporal information. For a given scalar time $\tau$, the $k$ sized time representation t2v is given by:
[…]
Radial Basis Functional Embedding
While Time2Vec incorporates the periodic properties of time, Radial Basis Functions (RBF) measure the similarity between a given scalar time and an agreed reference point. Taking inspiration from (Lewinson 2022), we propose a new type of embedding in which each element of the vector representation denotes the Gaussian-like distance between the scalar time and a trainable reference point, For a given scalar time $\tau$, the $k$-sized time representation is given by:
@@
Thus we have utlized two different methods of embedding scalar time. The RBF method is dependent upon the exponential distance between the time and a certain set of trainable points in space. On the other hand Time2Vec is a representation where the cyclic properties and emphasized. Together these two methods give complete information about the temporal properties of a sequence.
Our Model
Figure 2: The overall architecture of our model. For each session $S_{u}$, we send $n-1$ Action Sets, each having time $c_{u,i}^{(1)}$, date $c_{u,i}^{(2)}$, device $d_{u,i}^{(1)}$ and device control $d_{u,i}^{(2)}$ to get $\hat{y}_{u,k}$ as the output which is the prediction probabilities for each action
The model takes in a $N\times 10\times 4$ tensor as input and gives the predicted probabilities for all time intervals of a day. Currently, for each action, we have the time $c_{u,i}^{(1)}$, date $c_{u,i}^{(2)}$, device $d_{u,i}^{(1)}$ and device control $d_{u,i}^{(2)}$. Using this we first derive a new feature that contains the time difference between consecutive actions $c_{u,i}^{(3)}$ which denotes the time difference between $i^{th}$ and $(i-1)^{th}$ action in a sequence.
Embedding
Next, we rely on our embedding to convert the scalar values for different quantities into vectors. We rely on the linear embedding for the device and device control $e_{u,i}^{(1)}$ and $e_{u,i}^{(2)}$. For the temporal values, we utilize both the Time2Vec and Radial Basis Function approach to get a periodic embedding and a radial embedding respectively for the time and the date the action happened at giving us $z_{u,i}^{(1,1)}$, $z_{u,i}^{(1,2)}$, $z_{u,i}^{(2,1)}$, $z_{u,i}^{(2,2)}$. The time difference, however, is converted to an embedding using Time2Vec after multiplying it by a trainable factor as we are interested in the periodic properties itself giving us $z_{u,i}^{(3)}$.
Action Encoder
We make use of the famous Transformer Encoder Architecture (Vaswani et al. 2017) to encode all contextual information into a single vector. This includes the device, device control, and the two temporal embeddings of the day of action.
@@
Time Encoder
For each instance, we have $z_{u,i}^{(3)}$. First of all, we append a zero vector at the start, denoted as $z_{u,1}^{(3)}$ ensuring the sequence length is the same as that of the action encoder output. Now we pass this through a Temporal Convolution Network (Lea et al. 2017) to get $\hat{z}_{u,i}^{(3)}$ which contains the temporal information at the neighbourhood of each element of the sequence.
@@
Dataset and Experimental Setup
@@
SmartSense Dataset
The dataset (Jeon et al. 2022) contains real-world SmartThings data collected from various regions to evaluate performance. It comprises four log datasets and three routine datasets. The log datasets document device control histories executed by users of Bixby, serving as sequential instances for a general sequential recommendation.
The provided datasets are collected from SmartThings, a worldwide Internet of Things (IoT) platform with 62 million users. The log datasets are gathered over a month from four different countries: Korea (KR), USA (US), Spain (SP), and France (FR). Additionally, dictionaries mapping names to ids for various categories, such as day of weeks, hours, devices, controls, and device controls, are provided. The data is segregated by country into repositories, each corresponding to Korea, USA, Spain, and France, respectively.
For the log dataset, each instance is represented as a tensor of size $[N\times 10\times 5]$, where N denotes the number of instances. Each instance comprises ten consecutive actions performed by a Bixby service user, with each action represented as a vector of length 5, indicating the day of week, hour, device, control, and device control in order. The hour in a context is one of the 8-time ranges of 3 hours length: 0-3, 3-6, 6-9, 9-12, 12-15, 15-18, 18-21, and 21-24. The day, on the other hand, is an integer from 0 to 6.

### v2

@@ Abstract
The proliferation of IoT devices generates vast interaction data, offering insights into user behaviour. While prior work predicts what actions users perform, the timing of these actions—critical for enabling proactive and efficient smart systems—remains relatively underexplored. Addressing this gap, we focus on predicting the time of the next user action in smart environments. Due to the lack of public datasets with fine-grained timestamps suitable for this task and associated privacy concerns, we contribute a dataset of 11.6k sequences synthesized based on human annotations of interaction patterns, pairing actions with precise timestamps. To address, we introduce Timing-Matters, a Transformer-Encoder based method that predicts action timing, achieving 38.30% accuracy on the synthesized dataset, outperforming the best baseline by 6%, and showing 1–6% improvements on other open datasets. Our code and dataset will be publicaly released.
[…]
6795
1 Introduction
@@
In the era of technological advancement, the rapid growth of Internet of Things (IoT) solutions has resulted in an increase of smart devices within households, which is likely to almost double from 15.1 billion in 2020 to more than 29 billion IoT devices in 2030 [16]. This has changed the way we interact with technology in our daily lives. From smart thermostats to voice-activated assistants, these devices have seamlessly integrated into our homes, workplaces, and public spaces, providing convenience, efficiency, and connectivity like never before.
@@
Beyond task automation and real-time data provision, smart devices also have the potential to impact how machines understand and respond to human behaviour. These devices offer insights into our routines, lifestyle choices, and cognitive processes by continuously collecting data on user interactions, preferences, and habits. Leveraging these data can lead to the development of more intuitive and personalized technologies that anticipate the time of our needs and preferences, ultimately enhancing our over- all quality of life.
@@
A crucial yet often overlooked aspect of user behaviour prediction is the timing of actions. While existing research [8, 28, 29] focuses mainly on predicting the next user action, little attention has been given to forecasting when these actions will occur. Understanding and predicting the timing of user actions is crucial, as the majority of existing research on action prediction relies on time as a contextual feature; without accurate temporal modeling, predictions risk being misaligned with real user intent and behavior.
Consider a smart home system designed to help with daily routines. Suppose that the system can predict that a user will typically turn on the heater at 7:30 PM in winters. Even without applying automation to avoid energy wastage, the ability to recommend future events to users holds significant potential for smart devices and appliances. Other example (Figure 1) shows the activity of the user on a Sunday morning. As per daily/weekly behaviour the user is likely to start preparing for lunch. Without such temporal predictions, the system remains reactive, responding merely to user commands rather than anticipating needs.
Dataset provided in SmartSense [8] consists of real-world user interactions but lacks exact timestamps. To protect user privacy, Time is represented into 8 3-hour bins. To overcome this, we ask human annotators to simulate their behaviour in a smart home and create a new dataset with precise timestamps.
@@
By addressing this novel task of action time prediction, our research takes the first steps towards creating automated assistive AI systems capable of significantly improving the user experience in smart homes.
@@
Figure 1: Given the history of actions taken by the user we aim to predict the time of action
Our contributions are:
@@
    Synthesized Dataset: We introduce a novel dataset that includes data from 16 devices and 121 device controls. The dataset contains 11,665 instances, each representing a sequence of 10 consecutive actions with precise timestamps. (Section 5.1)
@@
    Timing-Matters: We propose a novel method, Timing-Matters, for predicting the timing of user actions. It achieves 38.30% accuracy on the synthesized dataset, outperforming the best baseline by approximately 6%, and shows consistent improvements of 1–6% across other datasets. (Section 4)
@@
    Context Window and Binning Analysis: We perform an extensive analysis on the impact of context window length and number of prediction bins when modeling this prediction problem. (Section 6.5)
  4. 4.
    Analysis of Regression vs. Classification Modeling: We conduct an in-depth study comparing the effectiveness of modeling this problem as a regression versus a classification task, highlighting the strengths and limitations of each approach. (Section 6.2)
2 Related Work
2.1 User Behaviour Prediction in Smart Homes
@@
The primary focus of research on user behaviour prediction in smart homes is on learning-based techniques, specifically deep learning and conventional machine learning techniques. Traditional methods like Hidden Markov Models (HMMs)  [15, 22] have been used to extract the user’s behaviour pattern and detect anomalies. However, HMM will fail due to the independence assumption [4] of HMM which makes it unable to consider context information. [1, 3, 5, 25, 32] exploit Long Short-Term Memory Networks (LSTM) to predict user behaviour in smart home scenarios. However, LSTM can only model long-term sequential influence [7], but misses out on the complex heterogeneous transitions and periodic sequential patterns caused by users’ routines and intents. SmartSense [8] adopts knowledge transfer to exploit user intent and employs a two-stage encoder to mine contextual information. However, it does not classify the routines before inputting the sequence into the encoder. Second, complex heterogeneous transitions are not modeled. Third, it does not take into account multilevel periodicity due to a lack of consideration of historical sequences. DeepUDI [28] employs the attention mechanism to consider the periodicity of human behaviour and the intent-aware encoder to consider intents. SmartUDI [29] considers routines, intent, and multilevel periodicity at the same time. However, none of the techniques attempts to predict the time when the user will perform the action.
2.2 Sequential User Behaviour Prediction
@@
Sequential user behaviour prediction methods can be categorized as follows. Traditional methods include Factorizing Personalized Markov Chains (FPMC) [20] factorize the users-locations matrix to generate user general preferences to complete the next location prediction. However, the location independence assumption prevents FPMC from capturing complicated sequential information [20]. CNN-based methods like Caser [24] employ CNN in both the time-axis and feature-axis to capture temporal dynamics in sequential recommendation. However, due to the limited receptive field of CNN, it cannot fully model long-term contextual information [11]. RNN-based methods like CARNN [18] and SIAR [19] incorporate contextual information into the RNN for sequential recommendation. However, CARNN and SIAR can not model multi-level periodicity because they only mine contextual information in a single sequence [18, 19]. Although DeepMove [6] uses attention-based RNN to model the correlation between the current sequence and historical sequences, it leads to suboptimal performance and long prediction time since it ambiguously considers all history sequences. SRGNN [27, 30] applies gated GNN to capture complex transition patterns among nodes for a session-based recommendation. However, SRGNN ignores the heterogeneity of the transition patterns caused by users’ intents. SASRec [9] utilizes unidirectional transformers to capture sequential patterns in sequences while considering the importance of correlations between behaviours. However, SASRec only captures the sequential patterns in single sequences, which prevents it from considering multi-level periodicity.
2.3 Time Prediction
Time prediction—forecasting the exact time a user will perform an action—has received limited attention in smart home research. Unlike time series forecasting, which predicts values over time, or event prediction, which focuses on what will happen, time prediction requires modelling user routines, contextual patterns, and multilevel periodicity. Traditional approaches, including regression and survival analysis [31], lack the capacity to capture sequential dependencies or dynamic contextual factors, making them unsuitable for behaviour-rich environments like smart homes. Classical point processes, such as Hawkes processes [12], model inter-event dependencies but assume stationary or parametric dynamics, which do not align with the complex, periodic nature of human activity [31, 2].Neural Temporal Point Processes (NTPPs) improve flexibility by learning event time distributions via RNNs or attention mechanisms [21]. However, NTPPs primarily model inter-event times rather than absolute time, and fail to incorporate periodicity, routines, or intent—core elements in smart home behaviour. Moreover, most surveyed models ignore contextual semantics (e.g., day of week, device interactions), limiting their applicability in domestic settings [2].
In summary, while existing event time models offer valuable insights, they fall short in modelling the structured, routine-driven, and context-rich temporal patterns necessary for accurate time prediction in smart homes.
3 Problem Statement
This paper addresses the problem of predicting the timing of a user’s next action within a smart environment, based on their sequence of preceding actions. Each historical action is represented by its device identifier, the specific control used, the day of the week, and the time it occurred. Our objective is to develop a model that leverages this sequence of past interactions to forecast *when* the subsequent action is most likely to take place, thereby providing valuable insights into temporal user behaviour patterns. For instance, given a sequence of n1n-1 user actions like the one illustrated in Fig. 1, the model’s task is to predict the time of the next ($n^{th}$) action.
@@ 3 Problem Statement
TIME OF ACTION PREDICTION
For each user instance $u$, we are given:
Input
A sequence of historical actions $S_{u}=[x_{u,1},\ldots,x_{u,(n-1)}]$, where each action $x_{u,i}=(d_{u,i}^{(1)},d_{u,i}^{(2)},c_{u,i}^{(1)},c_{u,i}^{(2)})$ is a quadruplet representing the indices of the device, device control, day of the week, and time, respectively, for the $i^{th}$ action in the sequence.
Output
The predicted time interval $\hat{y}_{u}$ in which the next ($n^{th}$) action is expected to occur. Our primary approach models this as a classification task, dividing the day into $96$ discrete 15-minute intervals. While this 96-class classification formulation is the main focus of our method description, we also evaluate the impact of different time granularities and compare the classification approach against regression-based modeling in our analysis (Section 6).
4 Methodology
4.0.1 Time2Vec Embedding
This is a widely used representation of time and ensures that both the sequential and cyclic patterns of time are captured [10]. This representation can be likened to a form of Fourier Transform. Except for the linear element, the rest of the elements split the time into different frequencies. This extracts the cyclic patterns present in the temporal information. For a given scalar time $\tau$, the $k$ sized time representation t2v is given by:
$\text{t2v}(\tau)[i]=\begin{cases}\omega_{i}\tau+\psi_{i}&\text{if }i=0\\ F(\omega_{i}\tau+\psi_{i})&\text{if }1\leq i\leq k-1\end{cases}$
Here, $\text{t2v}(t)[i]$ denotes the $i$th element of vector $\text{t2v}(t)$, $F$ is a periodic activation function with sine being the most common, and $\omega_{i}$ and $\psi_{i}$ are learnable parameters.
4.0.2 Radial Basis Functional Embedding
While Time2Vec incorporates the periodic properties of time, Radial Basis Functions (RBF) measure the similarity between a given scalar time and an agreed reference point. Taking inspiration from [14], we propose a new type of embedding in which each element of the vector representation denotes the Gaussian-like distance between the scalar time and a trainable reference point, For a given scalar time $\tau$, the $k$-sized time representation is given by:
@@ 4.0.2 Radial Basis Functional Embedding
Thus we have utilised two different methods of embedding scalar time. The RBF method is dependent upon the exponential distance between the time and a certain set of trainable points in space. On the other hand, Time2Vec is a representation where the cyclic properties and emphasized. Together these two methods give complete information about the temporal properties of a sequence.
4.1 Our Model
Figure 2: The overall architecture of our model. For each session $S_{u}$, we send n1n-1 Action Sets, each having time $c_{u,i}^{(1)}$, date $c_{u,i}^{(2)}$, time difference $c_{u,i}^{(3)}$, device $d_{u,i}^{(1)}$ and device control $d_{u,i}^{(2)}$ to get $\hat{y}_{u,k}$ as the output which is the prediction probabilities for each action
The model takes a $N\times 10\times 4$ tensor as input and gives the predicted probabilities for all time intervals of a day. Currently, for each action, we have the time $c_{u,i}^{(1)}$, date $c_{u,i}^{(2)}$, device $d_{u,i}^{(1)}$ and device control $d_{u,i}^{(2)}$. Using this, we first derive a new feature that contains the time difference between consecutive actions $c_{u,i}^{(3)}$ that denotes the time difference between $i^{th}$ and $(i-1)^{th}$ action in a sequence.
4.1.1 Embedding
Next, we rely on our embedding to convert the scalar values for different quantities into vectors. We rely on linear embedding for the device and device control $e_{u,i}^{(1)}$ and $e_{u,i}^{(2)}$. For the temporal values, we utilize both the Time2Vec and Radial Basis Function approach to get a periodic embedding and a radial embedding respectively for the time and the date the action happened at giving us $z_{u,i}^{(1,1)}$, $z_{u,i}^{(1,2)}$, $z_{u,i}^{(2,1)}$, $z_{u,i}^{(2,2)}$. The time difference, however, is converted to an embedding using Time2Vec after multiplying it by a trainable factor as we are interested in the periodic properties itself giving us $z_{u,i}^{(3)}$.
4.1.2 Action Encoder
We make use of the famous Transformer Encoder Architecture [26] to encode all contextual information into a single vector. This includes the device, device control, and the two temporal embeddings of the day of action.
@@ 4.1.2 Action Encoder
4.1.3 Time Encoder
For each instance, we have $z_{u,i}^{(3)}$. First of all, we append a zero vector at the start, denoted as $z_{u,1}^{(3)}$ ensuring the sequence length is the same as that of the action encoder output. Now we pass this through a Temporal Convolution Network [13] to get $\hat{z}_{u,i}^{(3)}$ which contains the temporal information at the neighbourhood of each element of the sequence.
@@ 4.1.5 Objective Function
5 Dataset and Experimental Setup
@@ 5.1 Dataset
5.1.1 SmartSense Dataset
The dataset [8] contains real-world SmartThings data collected from various regions to evaluate performance. It comprises four log datasets and three routine datasets. The log datasets document device control histories executed by users of Bixby, serving as sequential instances for a general sequential recommendation.
The provided datasets are collected from SmartThings, a worldwide Internet of Things (IoT) platform with 62 million users. The log datasets are gathered over a month from four different countries: Korea (KR), the USA (US), Spain (SP), and France (FR). Additionally, dictionaries mapping names to IDs for various categories, such as day of weeks, hours, devices, controls, and device controls, are provided.
For the log dataset, each instance is represented as a tensor of size $N\times 10\times 5]$, where N denotes the number of instances. Each instance comprises ten consecutive actions performed by a Bixby service user, with each action represented as a vector of length 5, indicating the day of week, hour, device, control ID, and device control in order. The hour in a context is one of the 8-time ranges of 3 hours length: 0-3, 3-6, 6-9, 9-12, 12-15, 15-18, 18-21, and 21-24. The day, on the other hand, is an integer from 0 to 6.
Dataset Task Precision(96) Precision(08) RMSE
FR R 0.1445 0.1445 63978.875
C 0.3567 0.3567 30225.904
AN R 0.2684 0.9329 8894.975
C 0.3865 0.7943 8804.339
Table 2: Comparison of R-Regression and C- Classification Methods